# The Disagreement Problem in Explainable Machine Learning: A Practitioner’s Perspective

### v5

@@ 1 Introduction
As machine learning (ML) models become integral in high-stakes domains like healthcare and finance, the interpretability of these models is crucial for ML practitioners and domain experts (e.g., doctors, and policymakers). Understanding model behavior is essential to identify errors and biases and gauge the reliability of model predictions, especially in complex models Doshi-Velez & Kim (2017). Several post hoc explanation methods have been proposed such as perturbation-based methods (e.g., LIME  Ribeiro et al. (2016b), SHAP Lundberg & Lee (2017)) and gradient-based methods (e.g., Gradient times Input  Simonyan et al. (2014), SmoothGrad Smilkov et al. (2017), Integrated Gradients Sundararajan et al. (2017), GradCAM Selvaraju et al. (2017)).
Prior research has investigated various aspects of explanation methods. For example, metrics have been developed to assess the faithfulness of explanations to the model’s behavior, such as comparing them to a model’s ground truth feature importances or using the "Remove and Retrain" (ROAR) approach Liu et al. (2021); Hooker et al. (2018). However, challenges arise due to the unavailability of ground truth in real-world applications and the impracticality of retraining models in certain settings Zhou et al. (2021). Studies have shown that popular explanation methods can produce inconsistent or misleading explanations, leading to risks like deploying biased models or making erroneous clinical decisions Aivodji et al. (2019); Slack et al. (2020).
However, while prior research has studied explanation methods, the (dis)agreement among different methods has not been extensively investigated. Practitioners often use multiple methods, but these can yield conflicting explanations. Therefore, resolving these disagreements is crucial to avoid reliance on misleading explanations with potentially severe consequences Slack et al. (2020). Yet, there is a lack of research on the extent to which explanations disagree and how practitioners resolve these disagreements.
In this work, we introduce and investigate the disagreement problem in explainable ML, a novel area of research that examines conflicts among post hoc explanation methods. We present the following key contributions:
@@ 2 Related Work
Our work is grounded in the extensive literature on explainable ML which we discuss in this section.
@@
Inherently Interpretable Models and Post Hoc Explanations. A variety of approaches have been developed to learn inherently interpretable models for tasks such as classification and clustering Letham et al. (2015); Lakkaraju et al. (2016); Bien & Tibshirani (2009); Kim et al. (2014); Lou et al. (2012); Caruana et al. (2015). However, complex models like deep neural networks often outperform these simpler models in terms of accuracy Ribeiro et al. (2016b). This has led to significant interest in post hoc explanation methods, which aim to understand the behavior of these complex models by approximating them using simpler models in a post hoc fashion. Post hoc explanation methods can be categorized into local explanation methods and global explanation methods. Local explanation methods, including LIME, SHAP, SmoothGrad, Integrated Gradients Ribeiro et al. (2016a); Lundberg & Lee (2017); Sundararajan et al. (2017); Smilkov et al. (2017) focus on explaining individual model predictions.
@@
Global explanation methods, on the other hand, aim to summarize the behavior of complex models as a whole Bastani et al. (2017); Lakkaraju et al. (2019). Our work is focused on analyzing the disagreements among explanations generated by the aforementioned local post hoc explanation methods.
@@
Analyzing and Evaluating Post hoc Explanations. Prior research has proposed several notions of explanation quality, such as faithfulness, stability, consistency, and sparsity Liu et al. (2021); Petsiuk et al. (2018); Slack et al. (2021); Zhou et al. (2021). Metrics have also been proposed to quantify these aspects of explanation quality Zhou et al. (2021); Carvalho et al. (2019); Gilpin et al. (2018); Liu et al. (2021). These properties and metrics have been leveraged to theoretically and empirically analyze the behavior of popular post hoc explanations Ghorbani et al. (2019); Slack et al. (2020); Dombrowski et al. (2019); Adebayo et al. (2018); Alvarez-Melis & Jaakkola (2018); Levine et al. (2019); Chalasani et al. (2020); Agarwal et al. (2021). The work that is closest to ours is the research by Neely et al. (2021) which demonstrates that certain post hoc explanation methods (e.g., LIME, Integrated Gradients, DeepLIFT, Grad-SHAP, Deep-SHAP, and attention based explanations) disagree with each other based on rank correlation (Kendall’s $\tau$) metric. However, their work neither formalizes the notion of explanation disagreement by leveraging practitioner inputs, nor studies how explanation disagreements are resolved in practice which are the key contributions of this work. This research has also inspired subsequent works that delve deeper into this topic Han et al. (2022); Banegas-Luna et al. (2023). For instance, Han et al. (2022) built on this work and theoretically analyzed multiple state-of-the-art explanation methods to understand the reasons behind their disagreements.
@@
Human Factors in Explainability. Many user studies have been conducted to evaluate how well humans can understand and use explanations Doshi-Velez & Kim (2017). Some of these studies have shown that data scientists often struggle to understand and effectively leverage state-of-the-art explanation techniques Kaur et al. (2020). Others have identified a variety of stakeholders across the model life cycle and highlighted the core goals of model explanations Hong et al. (2020); Chen et al. (2022). Prior works have also examined whether explanations help users in performing specific tasks Buçinca et al. (2021); Bansal et al. (2021); Fan et al. (2022); Sperrle et al. (2021); Chen et al. (2022); Vasconcelos et al. (2023); Fok & Weld (2023). However, none of these works focus on understanding the disagreement problem, the extent to which practitioners face it, and/or how they resolve it.
@@ 6 Discussion and Conclusion
The main takeaway from our work is the observation that state-of-the-art explanation methods consistently disagree with each other across various datasets and data modalities, as quantified by the metrics we outlined. Our empirical estimates of explanation disagreement closely align with human assessments of disagreements among leading explainable AI methods. Our focus was on highlighting the prevalence of disagreement within the literature on explainable AI and examining how practitioners address these disagreements. To this end, we also conducted extensive user studies (semi-structured interviews and an online user study) and empirical analyses. Our semi-structured interviews with machine learning practitioners highlighted that 84% of them regularly face explanation disagreements. While there exist metrics and frameworks to both evaluate and compare explanations output by state-of-the-art methods Agarwal et al. (2021; 2022); Han et al. (2022), several practitioners (86% of respondents in our online user study) admitted to using ad hoc heuristics or being uncertain about resolving explanation disagreements in practice. This highlights a critical gap between research and practice in explainable machine learning.
@@
In this work, we prioritized examining the prevalence of explanation disagreement rather than exploring its underlying causes, given the complexity of this analysis. Nonetheless, subsequent research building on our work has attempted to understand the sources of these disagreements. For example, Han et al. (2022) extended our work to demonstrate that state-of-the-art explanation methods all perform local linear approximations of underlying models but do so in different ways (e.g., using different loss functions and over different local neighborhoods). This can be a source of disagreement between the explanation methods. Furthermore, our results raise the question of how to resolve explanation disagreements when they occur. While this remains an active area of research, one approach to selecting among explanation methods is to first define the desired goal of an explanation and then choose the method that best meets this goal. For example, follow-up work by Han et al. (2022) defines local function approximation as the desired goal of an explanation, develops a theoretical framework to analyze how faithfully each method performs local function approximation and advises selecting the method that performs it most faithfully over a given input space. Given the high-stakes applications of model explanations, using a principled and systematic approach to select among explanation methods is preferable to the ad hoc strategies currently employed by practitioners. Finally, the extent of explanation disagreement would also depend on the specific metric being used to measure the disagreement. For instance, while our results broadly demonstrate a moderate to high degree of disagreement among all pairs of explanations across all data modalities, we observe that the extent of disagreement is slightly lower when we consider variants of the proposed metrics. See Appendix D.2 for additional details.
@@ Acknowledgements
This work is supported in part by the NSF awards IIS-2008461, IIS-2040989, IIS-2238714, and faculty research awards from Google, Adobe, JPMorgan, Harvard Data Science Initiative, and the Digital, Data, and Design (D^3) Institute at Harvard. The views expressed here are those of the authors and do not reflect the official policy or position of the funding agencies.

### v6

@@ 1 Introduction
As machine learning (ML) models become integral in high-stakes domains like healthcare and finance, the interpretability of these models is crucial for ML practitioners and domain experts (e.g., doctors and policymakers). Understanding model behavior is essential to identify errors and biases and gauge the reliability of model predictions, especially in complex models (Doshi-Velez & Kim, 2017). Several post hoc explanation methods have been proposed such as perturbation-based methods (e.g., LIME  (Ribeiro et al., 2016b), SHAP (Lundberg & Lee, 2017)) and gradient-based methods (e.g., Gradient times Input  (Simonyan et al., 2014), SmoothGrad (Smilkov et al., 2017), Integrated Gradients (Sundararajan et al., 2017), GradCAM (Selvaraju et al., 2017)).
Prior research has investigated various aspects of explanation methods. For example, metrics have been developed to assess the faithfulness of explanations to the model’s behavior, such as comparing them to a model’s ground truth feature importances or using the "Remove and Retrain" (ROAR) approach (Liu et al., 2021; Hooker et al., 2018). However, challenges arise due to the unavailability of ground truth in real-world applications and the impracticality of retraining models in certain settings (Zhou et al., 2021). Studies have shown that popular explanation methods can produce inconsistent or misleading explanations, leading to risks like deploying biased models or making erroneous clinical decisions (Aivodji et al., 2019; Slack et al., 2020).
However, while prior research has studied explanation methods, the (dis)agreement among different methods has not been extensively investigated. Practitioners often use multiple methods, but these can yield conflicting explanations. Therefore, resolving these disagreements is crucial to avoid reliance on misleading explanations with potentially severe consequences (Slack et al., 2020). Yet, there is a lack of research on the extent to which explanations disagree and how practitioners resolve these disagreements.
In this work, we formalize and investigate the disagreement problem in explainable ML, a novel area of research that examines conflicts among post hoc explanation methods. We present the following key contributions:
@@ 2 Related Work
Our work is grounded in the extensive literature on explainable ML, which we discuss in this section.
@@
Inherently Interpretable Models and Post Hoc Explanations. A variety of approaches have been developed to learn inherently interpretable models for tasks such as classification and clustering (Letham et al., 2015; Lakkaraju et al., 2016; Bien & Tibshirani, 2009; Kim et al., 2014; Lou et al., 2012; Caruana et al., 2015). However, complex models like deep neural networks often outperform these simpler models in terms of accuracy (Ribeiro et al., 2016b). This has led to significant interest in post hoc explanation methods, which aim to understand the behavior of these complex models by approximating them using simpler models in a post hoc fashion. Post hoc explanation methods can be categorized into local explanation methods and global explanation methods. Local explanation methods, including LIME, SHAP, SmoothGrad, and Integrated Gradients (Ribeiro et al., 2016a; Lundberg & Lee, 2017; Sundararajan et al., 2017; Smilkov et al., 2017) focus on explaining individual model predictions.
@@
Global explanation methods, on the other hand, aim to summarize the behavior of complex models as a whole (Bastani et al., 2017; Lakkaraju et al., 2019). Our work is focused on analyzing the disagreements among explanations generated by the aforementioned local post hoc explanation methods.
@@
Analyzing and Evaluating Post hoc Explanations. Prior research has proposed several notions of explanation quality, such as faithfulness, stability, consistency, and sparsity (Liu et al., 2021; Petsiuk et al., 2018; Slack et al., 2021; Zhou et al., 2021). Metrics have also been proposed to quantify these aspects of explanation quality (Zhou et al., 2021; Carvalho et al., 2019; Gilpin et al., 2018; Liu et al., 2021). These properties and metrics have been leveraged to theoretically and empirically analyze the behavior of popular post hoc explanations (Ghorbani et al., 2019; Slack et al., 2020; Dombrowski et al., 2019; Adebayo et al., 2018; Alvarez-Melis & Jaakkola, 2018; Levine et al., 2019; Chalasani et al., 2020; Agarwal et al., 2021). The work that is closest to ours is the research by Neely et al. (2021) which demonstrates that certain post hoc explanation methods (e.g., LIME, Integrated Gradients, DeepLIFT, Grad-SHAP, Deep-SHAP, and attention-based explanations) disagree with each other based on rank correlation (Kendall’s $\tau$) metric. However, their work neither formalizes the notion of explanation disagreement by leveraging practitioner inputs nor studies how explanation disagreements are resolved in practice, which are the key contributions of this work. This research has also inspired subsequent works that delve deeper into this topic (Han et al., 2022; Banegas-Luna et al., 2023). For instance, Han et al. (2022) built on this work and theoretically analyzed multiple state-of-the-art explanation methods to understand the reasons behind their disagreements.
@@
Human Factors in Explainability. Many user studies have been conducted to evaluate how well humans can understand and use explanations (Doshi-Velez & Kim, 2017). Some of these studies have shown that data scientists often struggle to understand and effectively leverage state-of-the-art explanation techniques (Kaur et al., 2020). Others have identified a variety of stakeholders across the model life cycle and highlighted the core goals of model explanations (Hong et al., 2020; Chen et al., 2022). Prior works have also examined whether explanations help users in performing specific tasks (Buçinca et al., 2021; Bansal et al., 2021; Fan et al., 2022; Sperrle et al., 2021; Chen et al., 2022; Vasconcelos et al., 2023; Fok & Weld, 2023). However, none of these works focus on understanding the disagreement problem, the extent to which practitioners face it, and/or how they resolve it.
@@ 6 Discussion and Conclusion
The main takeaway from our work is the observation that state-of-the-art explanation methods consistently disagree with each other across various datasets and data modalities, as quantified by the metrics we outlined. Our empirical estimates of explanation disagreement closely align with human assessments of disagreements among leading explainable AI methods. Our focus was on highlighting the prevalence of disagreement within the literature on explainable AI and examining how practitioners address these disagreements. To this end, we also conducted extensive user studies (semi-structured interviews and an online user study) and empirical analyses. Our semi-structured interviews with machine learning practitioners highlighted that 84% of them regularly face explanation disagreements. While there exist metrics and frameworks to both evaluate and compare explanations output by state-of-the-art methods (Agarwal et al., 2021; 2022; Han et al., 2022), several practitioners (86% of respondents in our online user study) admitted to using ad hoc heuristics or being uncertain about resolving explanation disagreements in practice. This highlights a critical gap between research and practice in explainable machine learning.
@@
In this work, we prioritized examining the prevalence of explanation disagreement rather than exploring its underlying causes, given the complexity of this analysis. Nonetheless, subsequent research building on our work has attempted to understand the sources of these disagreements. For example, (Han et al., 2022) extended our work to demonstrate that state-of-the-art explanation methods all perform local linear approximations of underlying models but do so in different ways (e.g., using different loss functions and over different local neighborhoods). This can be a source of disagreement between the explanation methods. Furthermore, our results raise the question of how to resolve explanation disagreements when they occur. While this remains an active area of research, one approach to selecting among explanation methods is to first define the desired goal of an explanation and then choose the method that best meets this goal. For example, follow-up work by Han et al. (2022) defines local function approximation as the desired goal of an explanation, develops a theoretical framework to analyze how faithfully each method performs local function approximation and advises selecting the method that performs it most faithfully over a given input space. Given the high-stakes applications of model explanations, using a principled and systematic approach to select among explanation methods is preferable to the ad hoc strategies currently employed by practitioners. Finally, the extent of explanation disagreement would also depend on the specific metric being used to measure the disagreement. For instance, while our results broadly demonstrate a moderate to high degree of disagreement among all pairs of explanations across all data modalities, we observe that the extent of disagreement is slightly lower when we consider variants of the proposed metrics. See Appendix D.2 for additional details.
@@ Acknowledgements
We thank the anonymous reviewers for their insightful comments and feedback on this paper. We also thank Javin Pombra for assisting with preliminary results during the initial phases of this study. Lastly, we are very grateful to Ana Lucic and Maurits J. R. Bleeker for pointing out the omission of a relevant study and helping us improve the contextualization of this work.